OpenStudy (anonymous):

Let f(x) = x2 – 1 and g(x) = 1 – 2x. Find the indicated values. f(g(–1)) and g(f(–1))

OpenStudy (anonymous):

\[g(-1)=?\]

OpenStudy (anonymous):

f(-1)?

OpenStudy (anonymous):

lol no \(g(-1)\) is a number

OpenStudy (anonymous):

\[g(x)=1-2x\\ g(-1)=1-2(-1)=?\]

OpenStudy (anonymous):

typo in my first question Let f(x) = x^2 – 1 and g(x) = 1 – 2x. Find the indicated values.

OpenStudy (anonymous):

there we go i forgot the ^ thingy

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

@satellite73

Nnesha (nnesha):

i guess he/she is not here but yes that is right f(-1) is 3

OpenStudy (anonymous):

satellite's gone way up to mars

Nnesha (nnesha):

nvm

OpenStudy (misty1212):

\[f(3)=3^2-1=9-1=8\]

OpenStudy (misty1212):

so \[f(g(-1))=f(3)=8\]

OpenStudy (anonymous):

what @misty1212 said

Nnesha (nnesha):

g**(-1)

OpenStudy (anonymous):

thanks @misty1212 and @Nnesha

Nnesha (nnesha):

wll np i guess